# Chip Placement with Diffusion Models

### v2

@@ 1 Introduction
By combining these ingredients, our method can generate placements for unseen netlists in a zero-shot manner, achieving results competitive with state-of-the-art on the ICCAD04 benchmark (Adya et al., 2004), also known as the IBM benchmark. Remarkably, our model accomplishes this without ever having trained on real-world circuit data.
@@ 2 Related Work
Google’s RL-based approach, CircuitTraining (Mirhoseini et al., 2021; Yue et al., 2022), employs a graph neural network (GNN) to generate netlist embeddings for multiple RL agents. While this method demonstrated state-of-the-art results, it relies on an initial placement generated using other methods (Cheng et al., 2023), and requires computationally expensive online training on new circuits. Several RL approaches follow to improve on runtime (Cheng & Yan, 2021; Lai et al., 2022; Gu et al., 2024), macro ordering (Chen et al., 2023), and proxy cost predictions (Zheng et al., 2023a, b; Wang et al., 2022; Ghose et al., 2021). Recently, ChiPFormer (Lai et al., 2023) demonstrated strong results on several benchmarks, combining offline and online RL to improve generalization. However their method still trains on in-distribution examples, and requires hours of online training on each new netlist for optimal results.
In contrast, Flora (Liu et al., 2022a) and GraphPlanner (Liu et al., 2022b) deviate from sequential placement formulations by leveraging a variational autoencoder (VAE) (Kingma & Welling, 2022) to generate placements. Flora further introduces a synthetic data generation scheme; however, it lacks variation in object sizes and restricts connections to only the nearest neighbors, which, as our experiments indicate, limits generalization to realistic circuit layouts (see Section 5.1). Furthermore, these generative models struggle to learn the underlying distribution of legal placements, frequently producing overlapping results.
Other approaches avoid the use of machine learning altogether. WireMask-BBO (Shi et al., 2023) utilizes black-box optimization algorithms to find optimal macro placements over continuous coordinates, while legalizing and evaluating the solution quality on a discrete grid. However, their usage of black-box optimization, such as evolutionary algorithms, leads to lengthy search times that must be started from scratch for each new circuit.
@@ 4.1 Generating Synthetic Data
Based on our results, we designed 2 synthetic datasets, v1 and v2, with parameters listed in Table 1.
@@ 4.4 Training and Evaluation
We evaluate the performance of our model on circuits in the publicly available ICCAD04 (Adya et al., 2004) (also known as IBM) benchmark. Because these circuits contain hundreds of thousands of small standard cells, we follow prior work (Mirhoseini et al., 2020) and cluster the standard cells into $512$ partitions using hMetis (Karypis et al., 1997). Each cluster is assigned to the nets of its constituent standard cells (nets within a single cluster are removed) with pins located at the cluster center, while the size of each cluster is the total area of its standard cells.
@@ 5.1.4 Zero-shot transfer to real circuits
We find that training on these broad-distribution datasets allow for effective zero-shot transfer to the real-world circuits in the IBM benchmark, with legality increasing significantly when training on v1 and v2, compared to the narrower v0 dataset. Although our simple data generation algorithm does not capture all the nuances of real circuits, such as the multimodal distributions of both edges and object sizes, it covers the important features well enough for our model to learn to produce reasonable placements on IBM circuits, some of which are shown in Figure 8.
@@ 5.3 Guided Sampling on Real Circuits
Table 4: Average HPWL and legality achieved by our models on the clustered IBM benchmark. DREAMPlace figures are included for comparison.
@@
Moreover, we find that our placements are significantly better than those produced by the state-of-the-art DREAMPlace (Lin et al., 2019), with $20\%$ lower HPWL on average. While we note that DREAMPlace, a mixed-size placer, is not optimized for placing clustered circuits, this result is nevertheless a strong indication of our method’s ability to generate high-quality placements for real circuits. This is especially remarkable, showing that models trained entirely on synthetic data can transfer effectively to real circuits in a zero-shot manner.
@@ 5.4 Mixed-Size Placement of Real Circuits
Table 5: Comparison of average HPWL ($10^{6}$) over 5 seeds using various techniques for mixed-size placement on the IBM benchmark.
@@
While our results (Table 4) have shown that our method can produce high-quality macro placements for clustered circuits, we also wish to investigate if these macro placements are useful for downstream tasks, particularly for mixed-size placement.
@@
To perform mixed-size placement (which requires placing all standard cells and macros) using our method, we first use our diffusion models to place clusters and macros (Figure 9(a)). We then initialize the position of each standard cell to the position of the corresponding cluster, and copy the positions of the macros (Figure 9(b)). Finally, we use DREAMPlace 4.1 (Lin et al., 2019) to place the standard cells, thus obtaining a full placement of standard cells and macros (Figure 9(c)). To elucidate the impact of our macro placements on the final placement quality, we keep the macro positions fixed during standard cell placement. This is in line with earlier works (Lai et al., 2022; Shi et al., 2023; Mirhoseini et al., 2020; Cheng & Yan, 2021) but contrasts with Lai et al. (2023), where the contribution of the macro placement technique is not as clear because the macros move significantly during standard cell placement.
In Table 5, we compare results from our method to other macro placement baselines, as well as DREAMPlace. The baselines include both the learning-based ChiPFormer (Lai et al., 2023) and MaskPlace (Lai et al., 2022), as well as the learning-free WireMask-BBO [^2] WireMask-BBO fails to place ibm12 and ibm15 so a lower-bound average is computed by substituting the smallest HPWL in the corresponding rows (Shi et al., 2023). All of these baselines use DREAMPlace for standard cell placement. We find that our method outperforms prior macro placement methods by a wide margin, while improving over DREAMPlace by $4\%$ on average. This indicates that the strong performance of our model on clustered macro placement transfers well to mixed-size placements. Moreover, our method, by design, can be easily applied zero-shot to new circuits and takes minutes to run, while other methods require RL fine-tuning or black box optimization, spending hours on each new circuit.

### v3

@@ 1 Introduction
By combining these ingredients, our method can generate placements for unseen netlists in a zero-shot manner, achieving results competitive with state-of-the-art on the ICCAD04 (or IBM) (Adya et al., 2004) and ISPD2005 (Nam et al., 2005b) benchmarks. Remarkably, our model accomplishes this without ever having trained on real circuit data.
@@ 2 Related Work
Google’s RL-based approach, CircuitTraining (Mirhoseini et al., 2021; Yue et al., 2022), employs a graph neural network (GNN) to generate netlist embeddings for multiple RL agents. While this method demonstrated state-of-the-art results, it requires computationally expensive online training on new circuits. Several RL approaches follow to improve on runtime (Cheng & Yan, 2021; Lai et al., 2022; Gu et al., 2024), macro ordering (Chen et al., 2023a), and proxy cost predictions (Zheng et al., 2023a, b; Wang et al., 2022; Ghose et al., 2021).
Several recent works have focused on reducing the long runtimes of RL methods. ChiPFormer (Lai et al., 2023) uses offline RL to reduce the amount of online training required. Despite strong results on several benchmarks, their method is constrained by the diversity and quantity of data available, requiring hours of online training on each new netlist for good results, even on in-distribution circuits. EfficientPlace (Geng et al., 2024) reduces the exploration needed by combining global tree search with a local policy. Although this improved efficiency, their method still relies on online training for the local policy, leading to runtimes of several hours per placement.
In contrast, Flora (Liu et al., 2022a) and GraphPlanner (Liu et al., 2022b) deviate from sequential placement formulations by leveraging a variational autoencoder (VAE) (Kingma & Welling, 2022) to generate placements. Flora further introduces a synthetic data generation scheme; however, it lacks variation in object sizes and restricts connections to only the nearest neighbors, which, as our experiments indicate, limits generalization to realistic circuit layouts (see Section 5.1 and Table 4). Furthermore, their models struggle to learn the underlying distribution of legal placements, frequently producing overlapping results.
Other approaches avoid the use of machine learning altogether. WireMask-BBO (Shi et al., 2023) utilizes black-box optimization algorithms to find optimal macro placements over continuous coordinates, while legalizing and evaluating the solution quality on a discrete grid. However, their usage of black-box optimization, such as evolutionary algorithms, leads to lengthy search times for each new circuit.
@@ 4.1 Generating Synthetic Data
Based on our results, we designed two synthetic datasets, v1 and v2, with parameters listed in Table 1.
@@ 4.4 Training and Evaluation
We evaluate the performance of our model on circuits in the publicly available ICCAD04 (Adya et al., 2004) (also known as IBM) and ISPD2005 (Nam et al., 2005a) benchmarks. Because these circuits contain hundreds of thousands of small standard cells, we follow prior work (Mirhoseini et al., 2020) and cluster the standard cells into $512$ partitions using hMetis (Karypis et al., 1997). Each cluster is assigned to the nets of its constituent standard cells (nets within a single cluster are removed) with pins located at the cluster center, while the size of each cluster is the total area of its standard cells.
@@ 5.1.4 Zero-shot transfer to real circuits
We find in Table 4 that training on these broad-distribution datasets allow for effective zero-shot transfer to the real-world circuits in the IBM benchmark, with legality increasing significantly when training on v1 and v2, compared to the narrower v0 dataset.
Moreover, models trained on our synthetic dataset far outperform those trained on the existing data generation algorithm proposed by Flora (Liu et al., 2022a, b). We see in Table 4 that our v1 dataset achieves a legality of $0.821$, far higher than Flora’s $0.264$, demonstrating that our dataset allows for much better zero-shot generalization.
Thus, although our simple data generation algorithm does not capture all the nuances of real circuits, such as the multimodal distributions of both edges and object sizes, it covers the important features well enough for our model to learn to produce reasonable placements on IBM circuits, some of which are shown in Figure 8.
@@ 5.3 Guided Sampling on Real Circuits
Table 4: Average HPWL and legality achieved by our models on the clustered IBM benchmark. DREAMPlace figures, and results for a model trained on Flora’s dataset are included for comparison.
@@
Moreover, we find that our placements are significantly better than those produced by the state-of-the-art DREAMPlace (Lin et al., 2019; Chen et al., 2023b), with $20\%$ lower HPWL on average. While we note that DREAMPlace, a mixed-size placer, is not optimized for placing clustered circuits, this result is nevertheless a strong indication of our method’s ability to generate high-quality placements for real circuits. This is especially remarkable, showing that models trained entirely on synthetic data can transfer effectively to real circuits in a zero-shot manner.
5.4 Macro-only Placement of Real Circuits
To compare our method to existing macro placement techniques, we also performed evaluations in the macro-only setting by removing all standard cells from the IBM and ISPD circuit netlists. For brevity, we present only averages for each benchmark in this section, with results for individual circuits found in Appendix B.
We see from Table 5 that our method outperforms existing macro placers by a large margin, reducing average HPWL by over $50\%$ compared to prior works in both the IBM and ISPD benchmarks.
Moreover, our method also outperforms prior works in terms of congestion, a commonly-used proxy for how difficult it is to route a given placement (Spindler & Johannes, 2007). Table 5 shows that compared to existing methods, our technique reduces average RUDY congestion by over $35\%$, measured using the method in Shi et al. (2023). The strong performance of our method on the congestion metric, despite not explicitly optimizing for it, is consistent with Shi et al. (2023)’s observation that RUDY congestion and HPWL are positively correlated.
Table 5: Average HPWL and RUDY congestion of our model compared to various baselines on the IBM and ISPD benchmarks in the macro-only setting. Our model significantly outperforms existing methods, achieving much lower HPWL and congestion.
Method HPWL ($10^{5}$) Congestion
IBM Average
MaskPlace $8.72$ $345$
WireMask-BBO $7.43$ $324$
ChiPFormer $7.33$ $336$
EfficientPlace $8.32$ $367$
Diffusion (Ours) $\mathbf{2.49}$ $\mathbf{196}$
ISPD Average
MaskPlace Timeout on $2/8$ circuits.
WireMask-BBO $154$ $1485$
ChiPFormer $116$ $836$
Diffusion (Ours) $\mathbf{45.9}$ $\mathbf{546}$
@@ 5.5 Mixed-Size Placement of Real Circuits
Table 6: Comparison of HPWL ($10^{6}$) averaged over 5 seeds, using various techniques for mixed-size placement, on the IBM benchmark.
@@
While our results (Table 4 and Table 5) have shown that our method can produce high-quality macro placements for clustered and macro-only circuits, we also wish to investigate if these macro placements are useful for downstream tasks, particularly for mixed-size placement.
@@
To perform mixed-size placement (which requires placing all standard cells and macros) using our method, we first use our diffusion models to place clusters and macros (Figure 9(a)). We then initialize the position of each standard cell to the position of the corresponding cluster, and copy the positions of the macros (Figure 9(b)). Finally, we use DREAMPlace 4.1 (Chen et al., 2023b) to place the standard cells, thus obtaining a full placement of standard cells and macros (Figure 9(c)). To elucidate the impact of our macro placements on the final placement quality, we keep the macro positions fixed during standard cell placement. This is in line with earlier works (Lai et al., 2022; Shi et al., 2023; Mirhoseini et al., 2020; Cheng & Yan, 2021) but contrasts with Lai et al. (2023), where the contribution of the macro placement technique is not as clear because the macros move significantly during standard cell placement.
In Table 6, we compare results from our method to other macro placement baselines, as well as DREAMPlace. The baselines include both the learning-based ChiPFormer (Lai et al., 2023) and MaskPlace (Lai et al., 2022), as well as the learning-free WireMask-BBO [^2] WireMask-BBO fails to place ibm12 and ibm15 so a lower-bound average is computed by substituting the smallest HPWL in the corresponding rows (Shi et al., 2023). All of these baselines use DREAMPlace for standard cell placement. We find that our method outperforms prior macro placement methods by a wide margin, while improving over DREAMPlace by $4\%$ on average. This indicates that the strong performance of our model on clustered macro placement transfers well to mixed-size placements. Moreover, our method, by design, can be easily applied zero-shot to new circuits and takes minutes to run (see Table 13), while other methods require RL fine-tuning or black box optimization, spending hours on each new circuit.
@@ Appendix A Implementation Details
Appendix B Additional Results
We present results for macro placements for each circuit in the IBM and ISPD benchmarks below. HPWL figures are shown in Table 9 and Table 10, congestion in Table 11 and Table 12, and runtime in Table 13 and Table 14.
In some cases, our values differ from those in the original papers because some baselines select and place only a subset of the macros, with differing selection criteria among baselines, while we place all macros for all baselines to ensure a fair comparison.
Table 9: Comparison of HPWL ($10^{5}$) for macro-only placements on the IBM benchmark. ibm05 has been omitted because it contains no macros.
Circuit MaskPlace WireMask-BBO ChiPFormer EfficientPlace Diffusion (Ours)
ibm01 $4.30$ $2.78$ $3.88$ $3.66$ $1.16$
ibm02 $5.54$ $4.19$ $5.05$ $4.42$ $2.68$
ibm03 $3.31$ $3.30$ $3.74$ $3.87$ $1.07$
ibm04 $6.91$ $5.43$ $5.96$ $6.10$ $2.40$
ibm06 $0.93$ $0.85$ $0.87$ $0.84$ $0.32$
ibm07 $2.67$ $2.66$ $2.36$ $3.42$ $0.78$
ibm08 $20.6$ $19.2$ $19.9$ $19.3$ $9.32$
ibm09 $2.45$ $1.76$ $1.77$ $2.57$ $0.44$
ibm10 $23.8$ $18.2$ $18.2$ $20.6$ $5.28$
ibm11 $4.15$ $3.75$ $3.25$ $4.70$ $0.78$
ibm12 $14.9$ $11.8$ $13.0$ $12.1$ $2.85$
ibm13 $4.58$ $4.41$ $4.02$ $5.37$ $1.05$
ibm14 $8.43$ $9.80$ $7.44$ $11.7$ $2.42$
ibm15 $4.68$ $7.77$ $2.67$ $5.98$ $1.06$
ibm16 $18.3$ $14.8$ $15.5$ $15.2$ $6.11$
ibm17 $16.8$ $12.2$ $13.7$ $17.9$ $3.20$
ibm18 $5.98$ $3.44$ $4.19$ $3.64$ $1.52$
Average $8.72$ $7.43$ $7.33$ $8.32$ $2.49$
Table 10: Comparison of HPWL ($10^{5}$) for macro-only placements on the ISPD benchmark.
Circuit MaskPlace WireMask-BBO ChiPFormer Diffusion (Ours)
adaptec1 $8.57$ $5.81$ $6.75$ $9.19$
adaptec2 $77.7$ $54.5$ $63.8$ $31.0$
adaptec3 $108$ $59.2$ $73.2$ $54.4$
adaptec4 $91.9$ $62.7$ $85.8$ $54.5$
bigblue1 $3.11$ $2.12$ $3.05$ $2.64$
bigblue2 Timeout $186$ $85.8$ $38.8$
bigblue3 $84.0$ $66.2$ $79.2$ $35.9$
bigblue4 Timeout $798$ $548$ $141$
Average $154$ $116$ $45.9$
Table 11: Comparison of congestion (RUDY estimator) for macro-only placements on the IBM benchmark. ibm05 has been omitted because it contains no macros.
Circuit MaskPlace WireMask-BBO ChiPFormer EfficientPlace Diffusion (Ours)
ibm01 $289$ $253$ $266$ $316$ $160$
ibm02 $228$ $243$ $205$ $257$ $178$
ibm03 $176$ $173$ $173$ $214$ $117$
ibm04 $449$ $483$ $490$ $480$ $260$
ibm06 $79.2$ $77.1$ $76.9$ $76.7$ $42.8$
ibm07 $154$ $164$ $160$ $177$ $83.0$
ibm08 $1232$ $1198$ $1261$ $1288$ $776$
ibm09 $127$ $119$ $111$ $153$ $49.1$
ibm10 $480$ $463$ $466$ $538$ $362$
ibm11 $180$ $183$ $172$ $240$ $69.2$
ibm12 $392$ $212$ $357$ $360$ $190$
ibm13 $163$ $202$ $177$ $209$ $86.0$
ibm14 $378$ $375$ $378$ $418$ $232$
ibm15 $162$ $173$ $173$ $227$ $69.2$
ibm16 $574$ $497$ $528$ $534$ $334$
ibm17 $531$ $464$ $488$ $483$ $204$
ibm18 $271$ $221$ $229$ $266$ $111$
Average $345$ $324$ $336$ $367$ $196$
Table 12: Comparison of congestion (RUDY estimator) for macro-only placements on the ISPD benchmark.
Circuit MaskPlace WireMask-BBO ChiPFormer Diffusion (Ours)
adaptec1 $312$ $139$ $140$ $149$
adaptec2 $1068$ $1084$ $1180$ $668$
adaptec3 $990$ $672$ $677$ $579$
adaptec4 $945$ $793$ $779$ $584$
bigblue1 $98.5$ $25.1$ $19.0$ $23.4$
bigblue2 Timeout $1924$ $500$ $523$
bigblue3 $970$ $955$ $956$ $391$
bigblue4 Timeout $6290$ $2436$ $1451$
Average $1485$ $836$ $546$
Table 13: Comparison of runtime (minutes) on the IBM benchmark. ibm05 has been omitted because it contains no macros.
Circuit MaskPlace WireMask-BBO ChiPFormer EfficientPlace DREAMPlace Diffusion (Ours)
ibm01 $154$ $209$ $98$ $54$ $0.308$ $1.85$
ibm02 $165$ $204$ $87$ $61$ $0.411$ $2.25$
ibm03 $123$ $217$ $75$ $61$ $0.393$ $2.17$
ibm04 $63$ $208$ $82$ $61$ $0.401$ $2.21$
ibm06 $34$ $224$ $80$ $29$ $0.229$ $2.38$
ibm07 $58$ $223$ $80$ $63$ $0.261$ $2.87$
ibm08 $75$ $207$ $105$ $79$ $0.260$ $3.38$
ibm09 $50$ $221$ $71$ $46$ $0.257$ $3.08$
ibm10 $516$ $228$ $236$ $268$ $0.455$ $4.50$
ibm11 $79$ $224$ $106$ $80$ $0.303$ $3.40$
ibm12 $390$ $253$ $196$ $206$ $0.469$ $5.24$
ibm13 $93$ $225$ $127$ $95$ $0.613$ $4.19$
ibm14 $393$ $266$ $187$ $216$ $0.760$ $5.95$
ibm15 $83$ $254$ $113$ $86$ $0.923$ $6.81$
ibm16 $107$ $217$ $137$ $130$ $0.784$ $7.88$
ibm17 $489$ $266$ $250$ $358$ $0.839$ $10.33$
ibm18 $60$ $216$ $93$ $58$ $0.742$ $8.06$
Average $172$ $227$ $124$ $114$ $0.475$ $4.39$
Table 14: Comparison of runtime (minutes) on the ISPD benchmark.
Circuit MaskPlace WireMask-BBO ChiPFormer DREAMPlace Diffusion (Ours)
adaptec1 $139$ $211$ $223$ $1.07$ $4.78$
adaptec2 $195$ $209$ $234$ $1.34$ $4.53$
adaptec3 $224$ $207$ $284$ $2.06$ $4.73$
adaptec4 $718$ $212$ $467$ $2.43$ $4.94$
bigblue1 $274$ $204$ $256$ $1.30$ $4.83$
bigblue2 Timeout $1396$ $5220$ $3.80$ $122$
bigblue3 $648$ $233$ $494$ $3.14$ $5.13$
bigblue4 Timeout $596$ $1210$ $8.86$ $18.9$
Average $408.5$ $1049$ $3.00$ $21.2$